### v2

@@ 5 Results
On the AMi-Br dataset, the best-performing model in terms of balanced accuracy was Virchow2 with LoRA finetuning, achieving an average balanced accuracy of 0.8135. While the highest AUROC of 0.9029 was obtained by the Swin Transformer model, Virchow2 closely followed with an AUROC of 0.9026, making it the most balanced performer overall. Other LoRA-fine-tuned models such as UNI (0.7952 mean balanced accuracy), Virchow (0.7878 mean balanced accuracy) and H-Optimus-0 (0.7846 mean balanced accuracy) also demonstrated strong performance, significantly improving over their respective linear probing counterparts. Among the end-to-end trained baseline models, the Swin Transformer achieved the highest balanced accuracy (0.8052) and AUROC (0.9029), outperforming both EfficientNetV2 and ViT. Foundation models with just linear probing lagged behind, with balanced accuracies ranging between 0.60-0.65.
On the AtNorM-Br dataset, the end-to-end trained ViT was the best performing model overall with a mean balanced accuracy of 0.7788 and AUROC 0.8710, while the Swin Transformer followed closely with a mean balanced accuracy of 0.7752 and an AUROC of 0.8683. Among the LoRA-fine-tuned foundation models, Virchow performed the best in terms of a mean balanced accuracy of 0.7696, followed by H-Optimus-1 (0.7658 mean balanced accuracy) and Virchow2 (0.7632 mean balanced accuracy) Other LoRA-fine-tuned models such as UNI2-h (0.7301 balanced accuracy), Prov-Gigapath (0.7263 balanced accuracy), UNI (0.7183 balanced accuracy) and H-Optimus-0 (0.7044 balanced accuracy) also showed consistent improvements over their linear probing variants. Foundation models with linear probing performed poorly, with balanced accuracies ranging between 0.59 and 0.65.
The AtNorM-MD dataset, which represents a distinct distribution shift, saw a general drop in performance across models. The end-to-end trained Swin Transformer performed the best, with a mean balanced accuracy of 0.7723 and an AUROC of 0.8806. Virchow pretrained with LoRA followed closely, with a balanced accuracy of 0.7705. Other LoRA-fine-tuned models also performed well, including Virchow2 (0.7424 balanced accuracy), H-Optimus-1 (0.7396 balanced accuracy), UNI (0.7069 balanced accuracy), Prov-Gigapath (0.7007 balanced accuracy), UNI2-h (0.6914 balanced accuracy) and H-Optimus-0 (0.6549 balanced accuracy), all of which substantially outperformed their linear probing counterparts. Among the end-to-end trained baselines, ViT followed closely after Swin Transformer, achieving 0.7494 balanced accuracy and 0.8720 AUROC. EfficientNetV2 achieved the least balanced accuracy of 0.6975 among the baselines. As with the other datasets, the foundation models without LoRA adaptation performed less robustly, with balanced accuracies ranging between 0.55 and 0.60.
@@ 6 Discussion
Our evaluation confirmed that all foundation models exhibited some degree of generalization to the AMF identification task, but the extent of their effectiveness varied considerably. The application of LoRA fine-tuning revealed particularly striking performance differences across models. Also, it is worthwhile to note that across nearly all datasets and performance metrics (balanced accuracy and AUROC), there consistently exists at least one end-to-end trained baseline model that outperforms both linear probing and LoRA-adapted foundation models.
We also notice a drop in performance for several settings, including both end-to-end trained models and foundation models (both linear probing and LoRAs-fine-tuned) in the two external test datasets, as reflected by the balanced accuracy and AUROC scores. We attribute this to the domain shift between the datasets. For the AtNorM-Br dataset, the observed performance may partially also be influenced by the reliance on a single expert involved in the classification of AMFs, potentially introducing systematic label biases and limiting generalizability, an effect that evens out when multiple raters are involved. Moreover, an image distribution shift cannot be ruled out, as the images were retrieved from different sources (labs, scanners, etc.), making the dataset more heterogeneous. For the AtNorM-MD dataset, where five annotators were involved, a systematic label bias is rather unlikely, however, the inclusion of multiple domains beyond breast cancer introduces a more pronounced image domain shift, encompassing a wider range of tissue types that pose significant challenges for model adaptation.
@@
CAB, VW, and CK acknowledge the support from the Austrian Research Fund (FWF, project number: I 6555). SB, TC, RK, and MA acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number: 520330054). KB acknowledges support by the German Research Foundation (DFG) project 460333672 CRC1540 EBM.

### v3

@@ 5 Results
On the AMi-Br dataset, the best-performing model in terms of balanced accuracy was Virchow2 with LoRA finetuning, achieving an average balanced accuracy of 0.8135 (sensitivity = 0.7511, specificity = 0.8760 of the atypical class). While the highest AUROC of 0.9029 was obtained by the Swin Transformer model, Virchow2 closely followed with an AUROC of 0.9026, making it the most balanced performer overall. Other LoRA-fine-tuned models such as UNI (0.7952 mean balanced accuracy), Virchow (0.7878 mean balanced accuracy) and H-Optimus-0 (0.7846 mean balanced accuracy) also demonstrated strong performance, significantly improving over their respective linear probing, as demonstrated by the pairwise Wilcoxon signed-rank tests for each test dataset (see Tab. 4). Among the end-to-end fine-tuned baseline models, the Swin Transformer achieved the highest balanced accuracy (0.8052) and AUROC (0.9029), outperforming both EfficientNetV2 and ViT. Foundation models with just linear probing lagged behind, with balanced accuracies ranging between 0.60-0.65.
On the AtNorM-Br dataset, the end-to-end fine-tuned ViT was the best performing model overall with a mean balanced accuracy of 0.7788 (sensitivity = 0.7109, specificity = 0.8466 of the atypical class) and AUROC 0.8710, while the Swin Transformer followed closely with a mean balanced accuracy of 0.7752 and an AUROC of 0.8683. Among the LoRA-fine-tuned foundation models, Virchow performed the best in terms of a mean balanced accuracy of 0.7696, followed by H-Optimus-1 (0.7658 mean balanced accuracy) and Virchow2 (0.7632 mean balanced accuracy) Other LoRA-fine-tuned models such as UNI2-h (0.7301 balanced accuracy), Prov-Gigapath (0.7263 balanced accuracy), UNI (0.7183 balanced accuracy) and H-Optimus-0 (0.7044 balanced accuracy) also showed consistent improvements over their linear probing variants. Foundation models with linear probing performed poorly, with balanced accuracies ranging between 0.59 and 0.65.
The AtNorM-MD dataset, which represents a distinct distribution shift, saw a general drop in performance across models. The end-to-end fine-tuned Swin Transformer performed the best, with a mean balanced accuracy of 0.7723 (sensitivity = 0.6548, specificity = 0.8897 of the atypical class) and an AUROC of 0.8806. Virchow pretrained with LoRA followed closely, with a balanced accuracy of 0.7705. Other LoRA-fine-tuned models also performed well, including Virchow2 (0.7424 balanced accuracy), H-Optimus-1 (0.7396 balanced accuracy), UNI (0.7069 balanced accuracy), Prov-Gigapath (0.7007 balanced accuracy), UNI2-h (0.6914 balanced accuracy) and H-Optimus-0 (0.6549 balanced accuracy), all of which substantially outperformed their linear probing counterparts. Among the end-to-end fine-tuned baselines, ViT followed closely after Swin Transformer, achieving 0.7494 balanced accuracy and 0.8720 AUROC. EfficientNetV2 achieved the least balanced accuracy of 0.6975 among the baselines. As with the other datasets, the foundation models without LoRA adaptation performed less robustly, with balanced accuracies ranging between 0.55 and 0.60.
@@ 6 Discussion
Our evaluation confirmed that all foundation models exhibited some degree of generalization to the AMF identification task, but the extent of their effectiveness varied considerably. The application of LoRA fine-tuning revealed particularly striking performance differences across models. Also, it is worthwhile to note that across nearly all datasets and performance metrics (balanced accuracy and AUROC), there consistently exists at least one end-to-end fine-tuned baseline model that outperforms both linear probing and LoRA-adapted foundation models.
We attribute this to not sufficiently discriminatory features extracted by the foundation models, of which many were not even trained on high magnification ($40\times$) patches (Zimmermann et al., 2024), leading to a considerable data distribution shift. Thus, the stronger adaptation of the model (by using rank-reduced adaptation through LoRA) provides considerable benefit for improving the extracted features, while using the full capacity of the model (full fine-tuning) yields even better results.
Most importantly, the primary scope of our study was not to identify a single best-performing model for the respective datasets for the task of AMF classification, but rather to compare different learning strategies for AMF classification - full fine-tuning of ImageNet-pretrained models, LoRA-based adaptation of foundation models, and linear probing of foundation models. To investigate this research question in a rigorous way, we performed statistical tests across all datasets to assess whether the performance differences between these learning strategies are statistically significant, and the resulting p-values are summarized in Table 4. These results show that the differences between linear probing and LoRA are statistically significant for all three datasets ($p\ll 0.001$), with LoRA consistently outperforming linear probing. Likewise, baseline models significantly outperform linear probing for all datasets ($p\ll 0.0001$).
We also note that the results of EfficientNet on the AMi-Br test dataset is better than reported in previous work (Bertram et al., 2025). This can be attributed to the fact that the latter used EfficientNet V2-S, while we use EfficientNet V2-M in our experiments.
We also notice a drop in performance for several settings, including both end-to-end fine-tuned models and foundation models (both linear probing and LoRAs-fine-tuned) in the two external test datasets, as reflected by the balanced accuracy and AUROC scores. We attribute this to the domain shift between the datasets. For the AtNorM-Br dataset, the observed performance may partially also be influenced by the reliance on a single expert involved in the classification of AMFs, potentially introducing systematic label biases and limiting generalizability, an effect that evens out when multiple raters are involved. Moreover, an image distribution shift cannot be ruled out, as the images were retrieved from different sources (labs, scanners, etc.), making the dataset more heterogeneous. For the AtNorM-MD dataset, where five annotators were involved, a systematic label bias is rather unlikely, however, the inclusion of multiple domains beyond breast cancer introduces a more pronounced image domain shift, encompassing a wider range of tissue types that pose serious challenges for model adaptation.
@@
Author contributions
SB wrote the main manuscript and carried out the experiments. VW, TAD, TC, RK, and CAB served as experts in the dataset annotation. SB and JA carried out the statistical analysis. CAB, KB, and MA co-wrote the manuscript and guided the project. RHJF organized and compiled the AtNorM-Br dataset with the support of CAB as pathology expert. CK provided scientific expertise. NP created Figure 1. All authors reviewed the manuscript.
@@
CAB, VW, and CK acknowledge the support from the Austrian Research Fund (FWF, project number: I 6555). SB, TC, RK, and MA acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number: 520330054). KB acknowledges support by the German Research Foundation (DFG) project 460333672 CRC1540 EBM. JA acknowledges support by the Bavarian State Ministry of the Sciences and the Arts (project FOKUS-TML). NP and MA acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number: 545049923)